The topic is find formula for the inverse of each of the following:
f(x)=x+5
f(x)+(x-2)square3 +5
g(x)=x+4dividedby3-x
f9x0+9x-2)square2+2
these are the four problems if you could please help solve them for me to better understand them. Thank you

Answer by stanbon(75887)

f(x)=x+5
Let's do the 1st one; the rest have so many typo problems it is
impossible to tell what you mean.
y=x+5
Interchange the x and the y to get:
x=y+5
Now, solve for y, as follows:
y=x-5
That is the inverse of y=x+5
Cheers,
Stan H.
Please edit or repost the other problems.